# Powerformer: A Transformer with Weighted Causal Attention for Time-series Forecasting

### v1

@@ 3.1.2 Incorporating Locality via Decaying Masks
We refer to this mechanism as WCMHA. By combining causal masking and smoothly decaying weights, WCMHA captures both the temporal ordering and localized dependencies inherent to many real-world processes.
3.2 Reweighting Functions for Locality
We explore multiple choices for the reweighting function $f(\Delta t)$:
  • Similarity Power Law: $f^{(\text{SPL})}(t)(\Delta t)=-(\Delta t)^{\alpha}.$
  • Weight Power Law: $f^{(\text{PL})}(t)(\Delta t)=-\alpha\,\log(\Delta t).$
  • Butterworth Filter: A frequency-based filter with a smooth transition resembling a step function, parameterized by its cutoff time $t_{c}$ and order $n$.
Figure 2: We show the weight power-law (solid line) and similarity power-law (dashed line) masks for varying $\alpha$. Panel (a) shows the contribution added to the attention scores and Panel (b) shows the subsequent effects on the attention weights after applying Softmax.
Each function has a tunable hyperparameter (e.g., $\alpha$ or $t_{c}$) that governs the decay rate of attention as $|\Delta t|$ grows. Figure 1c illustrates the conceptual shift in attention scores when applying these decays. Figures 2 and A1 illustrate the masks’ contributions to the attention score and weights for $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$, and $f^{(\text{BW})}_{n}(t)$, respectively. The $f^{(\text{PL})}(t)$ mask imparts a power-law envelope over the attention weight distribution, whereas the $f^{(\text{SPL})}(t)$ envelope is the exponential of a power-law. Therefore, $f^{(\text{SPL})}(t)$, which applies the power-law in the similarity score space, decays faster than $f^{(\text{PL})}(t)$ in weight space (as seen in Fig. 2).
3.3 Insights into WCMHA
Before integrating WCMHA into a state-of-the-art Transformer-based model, we investigate how causal and local masks reshape attention distributions in a vanilla Transformer. See Appendix A5.3 for details. Figure 3 compares attention score and weight distributions (encoder self-attention, decoder self-attention, and decoder cross-attention) on the Electricity dataset. We observe:
  • Encoder Self-Attention (modified by WCMHA) shifts toward higher positive values, emphasizing local contexts.
  • Decoder Cross-Attention (unchanged) retains a heavy-tailed distribution, indicating some long-range dependencies remain relevant.
  • Decoder Self-Attention (already causal) exhibits minimal distributional change.
These shifts consistently improve forecasting performance on real-world datasets (see Appendix A5.3). In many cases, WCMHA significantly outperforms standard MHA, especially for datasets with slowly decaying long-range dependencies. Interestingly, when WCMHA outperforms MHA it does so by much larger margins than when it underperforms. For example, on the ETTh2 dataset, WCMHA improves MSE and MAE by 16% and underperforms by 2%.
Figure 3: We show the attention score and weight distributions for both the benchmark Transformer (dotted black line) with MHA and our modified Transformer with WCMHA and $f^{(\text{PL})}(t)$ (solid colored lines). Panels (a), (b), and (c) correspond to the last encoder self-attention, decoder self-attention, and decoder cross-attention layers, respectively. The colored lines correspond to different mask decay times $(\alpha)$. These results are from the Electricity dataset with a 96 prediction length and 512 input length.
3.4 Powerformer: Locality, Causality, and Patching
Having shown that WCMHA provides a powerful mechanism for time-series, leveraging its full potential in practical scenarios requires an effective approach to represent temporal data. While incorporating WCMHA into modern Transformer time-series models, we wish to isolate the effects of WCMHA and its local and causal biases. To do so requires a simple architecture, and we focus on univariate time-series. Thus, we introduce Powerformer (see Fig. 1), a model that integrates Transformer, WCMHA, and PatchTST’s [27] successful patching framework.
Patching is a common tool in modern time-series models, and similar to $\mathbf{M}^{(\text{D})}$ it acts as a regularizer to remove high-frequency noise components. We decompose multivariate time-series into univariate channels before applying convolutional patching, similar to PatchTST [27]:
$\mathbf{X}^{(\text{Patched})}_{s,w}\;=\;\mathbf{X}*\mathbf{p}\Bigl{.}\Bigr{|}_{s},$
with patch length $w$ and convolution stride $s$. Patching captures local temporal contexts by converting segments of the raw series into compact embeddings for the Transformer.
Many state-of-the-art models only use the Transformer encoder [27, 49, 32] with a linear decoder head. Powerformer combines these insights with our previous Transformer experiments by using only the Transformer encoder with WCMHA and a similar linear readout head. Powerformer integrates two modern design choices while injecting causal masks and smoothly decaying attention weights, allowing it to better model nearby past time points.
Our design integrates two complementary strengths:
  • Locality and Causality: WCMHA introduces both temporal ordering and tailored decay, making it naturally suited for real-world time-series with power-law or exponential correlations.
  • Robust Patch Embeddings: Patching [27] efficiently captures short-term trends and reduces input length, mitigating quadratic complexity issues that can impede Transformers on long time-series.
Powerformer is expressive, computationally efficient, and sufficiently simple to interpret. Appendix A3.3 provides further details. Our empirical results confirm that imposing explicit locality and causality boosts forecasting accuracy and the interpretability of learned temporal dependencies.
4 Empirical Evaluation
We evaluate Powerformer [^1] https://github.com/khegazy/Powerformer, investigating the effects of WCMHA on the attention score and weight distributions, using the mean squared error (MSE) and mean absolute error (MAE) to quantify aggregate performance.
4.1 Datasets
@@
We evaluate Powerformer on 7 popular publicly-available real-world datasets [47, 42]: ETT (4 subsets), Weather, Electricity, and Traffic. We are particularly interested in the ETT subsets which measure the electricity transformer oil temperature at 2 different time scales (every 15 minutes and hourly) for two regions. Analyzing the same measurement in different locations and different time scales will illustrate how sensitive WCMHA is to sampling changes and distribution shifts. Section A2 provides a detailed description of these datasets, available at [42]. We note that Weather, Electricity, and Traffic have much larger timescales, spanning from one to multiple years (Table A1). These 7 datasets are large enough to capture slowly decaying temporal correlations, shown in Appendix A2. Illness [47] is traditionally used as a long-term forecasting benchmark. We omit this dataset as its temporal measurements (Table A1) and evaluation look-back window are more than an order of magnitude smaller than the other datasets. This short look-back window is typically the size of our filters.
4.2 Evaluation setup
@@
Following previous works [27, 42, 48, 39, 24, 32, 19, 47], we train for multiple prediction and input lengths. The prediction lengths are 96, 192, 336, and 720, while the input sequence lengths are 336 and 512. Each model is trained three times for each setting. Further details on our experimental process and hyperparameters are provided in Appendices A5.1 and A3.4, respectively.
@@
Table 1: We compare Powerformer’s performance on standard publicly-available time-series datasets to state-of-the-art Transformer models. The best results are bolded and second best are underlined. Powerformer and PatchTST are evaluated with look-back windows of 336 and 512, see Section A5.1 for selection criteria. TOTEM [32] and iTransformer [24] were evaluated with $T_{\text{seq}}=96$. All other models were evaluated with multiple look-back windows with the best results selected.
Powerformer PatchTST One-Fits-All TOTEM iTransformer FEDformer ETSformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.369 0.399 0.370 0.400 0.376 0.397 0.380 0.394 0.386 0.405 0.376 0.415 0.494 0.479
192 0.402 0.418 0.413 0.429 0.416 0.418 0.434 0.427 0.441 0.436 0.423 0.446 0.538 0.504
336 0.414 0.428 0.422 0.440 0.442 0.433 0.490 0.459 0.487 0.458 0.444 0.462 0.574 0.521
720 0.439 0.459 0.447 0.468 0.477 0.456 0.539 0.513 0.503 0.491 0.469 0.492 0.562 0.535
ETTh2 96 0.274 0.335 0.274 0.336 0.285 0.342 0.293 0.338 0.297 0.349 0.332 0.374 0.340 0.391
192 0.338 0.379 0.339 0.379 0.354 0.389 0.375 0.390 0.380 0.400 0.407 0.446 0.430 0.439
336 0.325 0.379 0.331 0.380 0.373 0.407 0.422 0.431 0.428 0.432 0.400 0.447 0.485 0.479
720 0.376 0.419 0.379 0.422 0.406 0.441 0.610 0.567 0.427 0.445 0.412 0.469 0.500 0.497
ETTm1 96 0.290 0.342 0.290 0.342 0.292 0.346 0.320 0.347 0.334 0.368 0.326 0.390 0.375 0.398
192 0.329 0.369 0.332 0.369 0.332 0.372 0.379 0.382 0.377 0.391 0.365 0.415 0.408 0.410
336 0.359 0.389 0.366 0.392 0.366 0.394 0.406 0.402 0.426 0.420 0.392 0.425 0.435 0.428
720 0.412 0.421 0.420 0.424 0.417 0.421 0.471 0.438 0.491 0.459 0.446 0.458 0.499 0.462
ETTm2 96 0.162 0.252 0.165 0.255 0.173 0.262 0.176 0.253 0.180 0.264 0.180 0.271 0.189 0.280
192 0.222 0.292 0.220 0.292 0.229 0.301 0.247 0.302 0.250 0.309 0.252 0.318 0.253 0.319
336 0.275 0.327 0.278 0.329 0.286 0.341 0.317 0.348 0.311 0.348 0.324 0.364 0.314 0.357
720 0.351 0.380 0.367 0.385 0.378 0.401 0.426 0.410 0.412 0.407 0.410 0.420 0.414 0.413
Weather 96 0.147 0.198 0.149 0.198 0.162 0.212 0.165 0.208 0.174 0.214 0.238 0.314 0.197 0.281
192 0.191 0.239 0.194 0.241 0.204 0.248 0.207 0.250 0.221 0.254 0.275 0.329 0.237 0.312
336 0.243 0.279 0.245 0.282 0.254 0.286 0.257 0.291 0.278 0.296 0.339 0.377 0.298 0.353
720 0.310 0.329 0.314 0.334 0.326 0.337 0.326 0.340 0.358 0.347 0.389 0.409 0.352 0.388
Electricity 96 0.129 0.223 0.129 0.222 0.139 0.238 0.178 0.263 0.148 0.240 0.186 0.302 0.187 0.304
192 0.145 0.240 0.147 0.240 0.153 0.251 0.187 0.272 0.162 0.253 0.197 0.311 0.199 0.315
336 0.162 0.257 0.163 0.259 0.169 0.266 0.199 0.285 0.178 0.269 0.213 0.328 0.212 0.329
720 0.198 0.289 0.197 0.290 0.206 0.297 0.236 0.318 0.225 0.317 0.233 0.344 0.233 0.345
Traffic 96 0.365 0.251 0.360 0.249 0.388 0.282 0.523 0.303 0.395 0.268 0.376 0.415 0.607 0.392
192 0.382 0.258 0.379 0.256 0.407 0.290 0.530 0.303 0.417 0.276 0.423 0.446 0.621 0.399
336 0.391 0.264 0.392 0.264 0.412 0.294 0.549 0.311 0.433 0.283 0.444 0.462 0.622 0.396
720 0.430 0.285 0.432 0.286 0.450 0.312 0.598 0.331 0.467 0.302 0.469 0.492 0.632 0.396
$1^{\text{st}}/2^{\text{nd}}$ 47 / 8 17 / 33 3 / 13 1 / 3 0 / 1 0 / 0 0 / 0
4.3 Empirical Results

### v2

@@ 3.2 Recency-Biased Causal Attention
By combining causal masking and smoothly decaying weights, RBCA captures temporal ordering and localized dependencies inherent to real-world processes.
3.3 Recency Bias Functions
Generally, previous NLP works consider sliding windows (Beltagy et al., 2020) and exponentially decaying recency biases (ALiBi (Press et al., 2022)). For RBCA, these are given by
$f^{(\text{BW})}_{n}(t)=\left(1+\left(\frac{z(t)}{t_{c}}\right)^{2n}\right)^{-1/2},\quad f^{(\text{E})}(t)=-\alpha t,$
respectively, where $\alpha$ is the decay time constant, and $f^{(\text{BW})}_{n}(t)$ is the Butterworth filter that mimics a smoothed step function (Figs. 1c and 2). The Butterworth filter (Section S1.2) depends on the cutoff time $t_{c}$, the decay order $n$, and the digital filter gain $z(t)$.
Here, we leverage the fact that pairwise correlations in many physical systems follow a power-law distribution (Clauset et al., 2009). We note that attention weights are analogous to pairwise correlations, as they measure a pairwise similarity. Consequently, we make two power-law recency biases
$f^{(\text{PL})}(t)=-\alpha\,\log(t),\quad f^{(\text{SPL})}(t)=-(t)^{-\alpha}.$
$f^{(\text{PL})}(t)$ induces a power-law envelope onto the attention weights, and $f^{(\text{SPL})}(t)$ adds a power-law bias to the attention score that results in an exponential of a power-law after the softmax.
3.4 Regularization and Reduced Complexity
The recency bias is a regularizer on the input data, as it limits MHA’s ability to attend to inputs far away. Intuitively, the recency bias acts as a hard-constraint that forces the method to attend to recent measurements. Some recency biases, depending on the decay rate $\alpha$, sufficiently attenuate the weights to 0 at some time $\tau$. Figures 1c and 2 show such instances. In these cases, the attention complexity can be reduced from quadratic in input length $T$ to linear: O($\tau T$).
3.5 Interpreting Power-Law Biases
Power-law recency biases strike a balance between sampling sinusoidal and slowly developing signals over long periods of time with sampling nearby quickly varying signals. To develop this intuition, we consider two limits of signal types: periodic systems (which benefit from long-time windows) and systems with fast exponentially decaying pairwise correlations (which benefit from a quickly decaying recency bias). Sinusoidal signals benefit from long and equally weighted windows, while short exponentially decaying signals benefit from swiftly decaying biases. The power-law biases are a combination of the benefits from the exponential bias and the long-time window. At recent times, the mask decays the fastest: analogous to the exponential decay that amplifies the most recent information. At longer times, the heavy tail allows the model to sample many more periods of an oscillation than other biases (e.g. exponential). The heavy tail filters for long-term signals since higher-frequency oscillations and fast variations will be averaged out, but long-term oscillations and slowly developing signals on the order of the tail’s length will remain. By attending to attenuated distant tokens, the power law enables RBCA to capture (or remember) long-term signals, providing a more expressive class of inductive biases that outperform exponential decay in practice.
4 EXPERIMENTS
We investigate the general effects of recency biases (RBCA) on Transformers for causal and local tasks. This includes both NLP and time-series tasks. First, we test if the recency bias endows Transformer with capabilities similar to RNN’s read, ignore, and write. We see that some biases do give Transformer these capabilities. Second, we build upon this insight by applying RBCA to the original encoder-decoder Transformer and testing on standard time-series tasks. Lastly, we continue investigating the benefits of RBCA by applying it to a simple state-of-the-art time-series model (PatchTST (Nie et al., 2023)) and find that it further improves performance and outperforms other models that claim to beat PatchTST.
4.1 Addressing Transformer Glitching
Prior work (Liu et al., 2023) showed that Transformers struggle to reliably perform simple, yet important, RNN tasks together: write information to a state, ignore input information, and read from the stored state. These tasks are important for filtering and recalling important information in long sequences. RNNs can easily perform this task, likely due to their causal structure and implicit recency bias. Here, we investigate whether the causal and recency biases from RBCA endow Transformer with similar capabilities.
Evaluation
To test these capabilities, (Liu et al., 2023) developed the flip-flop test: a string of alternating characters $c\in\{i,w,r\}$ and numbers $n\in\{0,1\}$. Each character instructs the model to ignore ($i$) the following number, write ($w$) the following number to memory, or read ($r$) the latest number that was written. For the example $w0i1i0w1i0r$, the model reads $1$ since it was the last written number. The out-of-distribution (OOD) test contains many ignores ($i$) after the last write ($w$) and tests Transformer’s ability to generalize its read, ignore, and write capabilities to longer sequences. We train and test a 2-layer encoder-only Transformer with and without RBCA. Each attention head has a different $\alpha$. Section S5.1 provides details on the model, evaluation, and results.
Attention Analysis
We first investigate the effects of RBCA on the attention weight distributions of ALiBi (one of the most prominent NLP recency biases) and our physically-inspired power-law bias: $f^{(\text{E})}(t)$ and $f^{(\text{PL})}(t)$ respectively. Figure 3 shows these attention weight distributions for the best attention heads, full results are in Section S5.1. Most importantly, the attention weight between the ‘read’ and correct number pair needs to be high, while the weight between the ‘read’ and the incorrect numbers should be close to 0. For ALiBi ($f^{(\text{E})}(t)$), the amount of weights equal to 0 and 1 for the correct number pair are similar (Fig. 3a). Conversely, for $f^{(\text{PL})}(t)$, all weights corresponding to the correct number are nonzero, with almost 90% equal to 1 (Fig. 3c). When identifying the correct ‘write’ command, almost all the weights corresponding to the correct ‘write’ are 0 for ALiBi (Fig. 3b), indicating that it is generally unable to find it. For $f^{(\text{PL})}(t)$, we see a distinct shift between the correct and incorrect ‘write’, with all weights greater than 0 for the correct ‘read’ and ‘write’ pair: (Fig. 3d). After observing these fundamental changes to the attention weight structure, we now explore these biases’ performances.
Figure 3: The attention weights corresponding to the flip flop ‘read’ command and the correct/incorrect ‘write’ command and number. We show these distributions for the ALiBi ($f^{(\text{E})}(t)$) and the $f^{(\text{PL})}(t)$ biases.
Results
When comparing the RBCA accuracy on the OOD test set against standard MHA, both power-law biases ($f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$) consistently outperform, while both NLP biases ($f^{(\text{E})}(t)$ and $f^{(\text{BW})}_{n}(t)$) underperform. As noted by (Press et al., 2022), the OOD accuracy varies during training, but $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ consistently hover around 100%, half the time being exactly 100%. Both NLP-inspired recency biases result in worse accuracy compared to vanilla MHA. MHA hovers around 85%, ALiBi ($f^{(\text{E})}(t)$) hovers around 75%, and sliding window ($f^{(\text{BW})}_{n}(t)$) performs very poorly. Section S5.1 provides full OOD test results. These results indicate that recency biases significantly affect sequential read, write, and ignore capabilities. The correct recency bias can be sufficient to endow the Transformer with these capabilities.
Figure 4: After training Powerformer, we show the attention weights without the mask $\mathbf{M}^{(\text{C,L})}$ (a and b) and with the mask (c and d) for the first (a and c) and last (b and d) encoder layers. The linear trend illustrates the power law. These results are from the Weather dataset with $f^{(\text{PL})}(t)$, $\alpha=1.0$, and input/prediction lengths 512/96.
4.2 RBCA in Vanilla Transformer
Given the success of power-law-biased RBCA on flip-flop, we further evaluate RBCA’s capabilities on more challenging causal and local tasks, namely time-series forecasting. Here, we take a principled approach by first testing the effects of RBCA on a standard encoder-decoder Transformer. This is important, because adding more features to the model will convolute RBCA’s effects and potentially confuse the results.
Datasets
@@ 4.2 RBCA in Vanilla Transformer
Our time-series tasks use 7 standard benchmark datasets (Zhou et al., 2021; Wu et al., 2021): ETT (4 subsets), Weather, Electricity, and Traffic. The ETT subsets measure the electricity transformer oil temperature at 2 different time scales (every 15 minutes and hourly) for two regions. These are particularly interesting for testing the RBCA’s sensitivity to sampling changes and distribution shifts. Section S2 provides a detailed description of these datasets, available in (Wu et al., 2021). We note that Weather, Electricity, and Traffic have much larger timescales, spanning from one to multiple years (Table S1). These 7 datasets are large enough to capture slowly decaying temporal correlations (Appendix S2).
Evaluation
@@
Following previous works (Nie et al., 2023; Wu et al., 2021; Zhou et al., 2022; Woo et al., 2022; Liu et al., 2024; Talukder et al., 2024; Kitaev et al., 2020; Zhou et al., 2021), we train for multiple prediction lengths (96, 192, 336, and 720) and input lengths (336 and 512). Each model is trained three times for each setting with different seeds. For each dataset, we treat the recency bias, decay rate, and input sequence length as hyperparameters. We select one input sequence for the entire dataset, and a recency bias and decay rate for each prediction length. Sections S5.2 and S3.4 provide details on our experimental process and hyperparameters, respectively. All benchmarks are trained with varying input sequences, and the optimal model is selected in the same way.
For each dataset, we evaluate the $f^{(\text{PL})}(t)$ and $f^{(\text{SPL})}(t)$ biases. We evaluate $f^{(\text{BW})}_{n}(t)$ on the ETT and weather datasets only, due to its poor performance on these datasets and the Transformer glitching benchmark. We do not evaluate $f^{(\text{E})}(t)$ due to its similarity to $f^{(\text{SPL})}(t)$ and inferior performance on the Transformer glitching benchmark.
Results
In the vast majority of cases, RBCA outperforms standard MHA, especially for datasets with slowly decaying long-range dependencies. When RBCA outperforms MHA, it does so by much larger margins than when it underperforms. For example, on the ETTh2 dataset RBCA and MHA outperform each other the same number of times, however, RBCA improves MSE and MAE by 16% and underperforms by 2%. Table S4 shows the full MSE and MAE results.
4.3 Powerformer
To further investigate RBCA’s effects on time-series forecasting, we add it to modern time-series architectures to determine how RBCA may affect state-of-the-art models. Similar to our previous tests, our goal is to understand the effects of recency biases on attention. For this reason, we develop a simple time-series model, allowing us to isolate the recency bias effects. Adding many features convolutes the recency bias’ effects and confuses the source of the outcome.
Model
We introduce Powerformer, a time-series Transformer variant built atop PatchTST (Nie et al., 2023) by replacing MHA with RBCA. Powerformer integrates the encoder-only architecture, RBCA, and PatchTST’s (Nie et al., 2023) successful patching framework. PatchTST follows common time-series architecture designs: an encoder-only Transformer with a linear readout head. We select PatchTST because it does not alter the Transformer, but instead only makes changes to the input embeddings by patching. The lack of alterations to the attention mechanism in PatchTST will isolate the effects of RBCA on both attention weight distributions and forecasting performance. While there have been more advances in time-series forecasting since PatchTST, such advancements make larger changes to the Transformer and attention architecture. We believe that many of these changes will potentially confuse the results by making it challenging to determine where the improvement came from. Appendix S3.3 provides further details on Powerformer’s architecture.
Evaluation
We evaluate Powerformer on the same datasets and in the same manner as done in Section 4.2. Sections S5.2 and S3.4 provide further details on our experimental process and hyperparameters, respectively. We benchmark Powerformer against models that focus on modifying the attention mechanism and models that use recency biases. These models take a similar approach to Powerformer and are thus a fair comparison. We do not compare against all state-of-the-art models, as they contain techniques Powerformer purposely lacks, but can be added to Powerformer, or RBCA could be added to other models. Either expanding Powerformer or adding RBCA to other models is an exciting direction for future work.
Attention Analysis
Interestingly, Powerformer strengthens our causal and recency biases beyond $\mathbf{M}^{(\text{C,R})}$, which is a key insight into the importance of the causal and recency biases. At first glance, this is counterintuitive because the causal and recency biases act as regularizers that remove (and diminish contributions from) parts of the input sequence. The removed data could otherwise be used to help the model memorize the output and overfit. Moreover, since the recency bias mask ($\mathbf{M}^{(\text{R})}$) is constant, Powerformer has to capability to mitigate it by learning to counter the mask’s regularization effects.
Rather than learning to remove the recency bias regularization, Powerformer strengthens the recency bias. To demonstrate this behavior, we look at the attention weight distribution with and without $\mathbf{M}^{(\text{C,R})}$on Powerformer trained with $f^{(\text{PL})}(t)$ ($\alpha=1$) and $\mathbf{M}^{(\text{C,R})}$. Figure 4 shows the attention weight distribution as a function of time delay between embeddings. The first encoder layer without the mask (Fig. 4a) is relatively invariant to causality and locality (unbiased). When $\mathbf{M}^{(\text{C,R})}$ is applied to the first layer (Fig. 4c) the non-causal weights vanish and the power-law structure emerges. This behavior is expected, unlike the behavior in the last layer. In the last layer, the attention weight distribution without $\mathbf{M}^{(\text{C,R})}$ (Fig. 4b) is entirely causal with a strong recency bias. That is, even without applying the causal and recency bias, Powerformer is enforcing causality and the power-law recency bias, rather than trying to remove this regularization. Therefore, once $\mathbf{M}^{(\text{C,R})}$ is applied to the last layer (Fig. 4d), the recency bias is further increased, as seen by the steeper slope. Therefore, Powerformer learns embeddings that further strengthen the effects of the recency bias, which is strong evidence that these biases are crucial in time-series forecasting.
@@ 4.3 Powerformer
Table 1: Powerformer’s performance on standard publicly-available time-series datasets compared to similar state-of-the-art Transformer models. The best and second best results are bolded and underlined, respectively.
Powerformer PatchTST iTransformer One-Fits-All FEDformer ETSformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.361 0.390 0.370 0.400 0.401 0.417 0.376 0.397 0.376 0.415 0.494 0.479
192 0.395 0.410 0.413 0.429 0.440 0.445 0.416 0.418 0.423 0.446 0.538 0.504
336 0.406 0.420 0.422 0.440 0.455 0.454 0.442 0.433 0.444 0.462 0.574 0.521
720 0.434 0.455 0.447 0.468 0.520 0.511 0.477 0.456 0.469 0.492 0.562 0.535
ETTh2 96 0.269 0.334 0.274 0.336 0.302 0.358 0.285 0.342 0.332 0.374 0.340 0.391
192 0.330 0.375 0.339 0.379 0.369 0.400 0.354 0.389 0.407 0.446 0.430 0.439
336 0.323 0.380 0.331 0.380 0.407 0.430 0.373 0.407 0.400 0.447 0.485 0.479
720 0.370 0.417 0.379 0.422 0.436 0.462 0.406 0.441 0.412 0.469 0.500 0.497
ETTm1 96 0.285 0.342 0.290 0.342 0.301 0.355 0.292 0.346 0.326 0.390 0.375 0.398
192 0.328 0.368 0.332 0.369 0.342 0.380 0.332 0.372 0.365 0.415 0.408 0.410
336 0.357 0.385 0.366 0.392 0.376 0.401 0.366 0.394 0.392 0.425 0.435 0.428
720 0.410 0.414 0.420 0.424 0.440 0.438 0.417 0.421 0.446 0.458 0.499 0.462
ETTm2 96 0.163 0.251 0.165 0.255 0.179 0.271 0.173 0.262 0.180 0.271 0.189 0.280
192 0.218 0.290 0.220 0.292 0.236 0.310 0.229 0.301 0.252 0.318 0.253 0.319
336 0.273 0.326 0.278 0.329 0.284 0.342 0.286 0.341 0.324 0.364 0.314 0.357
720 0.358 0.380 0.367 0.385 0.370 0.394 0.378 0.401 0.410 0.420 0.414 0.413
Weather 96 0.147 0.197 0.149 0.198 0.163 0.211 0.162 0.212 0.238 0.314 0.197 0.281
192 0.191 0.239 0.194 0.241 0.204 0.253 0.204 0.248 0.275 0.329 0.237 0.312
336 0.243 0.279 0.245 0.282 0.256 0.291 0.254 0.286 0.339 0.377 0.298 0.353
720 0.310 0.329 0.314 0.334 0.326 0.337 0.326 0.337 0.389 0.409 0.352 0.388
Electricity 96 0.129 0.223 0.129 0.222 0.131 0.226 0.139 0.238 0.186 0.302 0.187 0.304
192 0.145 0.240 0.147 0.240 0.152 0.247 0.153 0.251 0.197 0.311 0.199 0.315
336 0.162 0.257 0.163 0.259 0.168 0.263 0.169 0.266 0.213 0.328 0.212 0.329
720 0.198 0.289 0.197 0.290 0.191 0.284 0.206 0.297 0.233 0.344 0.233 0.345
Traffic 96 0.365 0.252 0.360 0.249 0.352 0.258 0.388 0.282 0.576 0.359 0.607 0.392
192 0.382 0.258 0.379 0.256 0.368 0.267 0.407 0.290 0.610 0.380 0.621 0.399
336 0.391 0.265 0.392 0.264 0.384 0.273 0.412 0.294 0.608 0.375 0.622 0.396
720 0.430 0.286 0.432 0.286 0.417 0.290 0.450 0.312 0.621 0.375 0.632 0.396
$1^{\text{st}}/2^{\text{nd}}$ 46 / 7 9 / 38 6 / 3 0 / 10 0 / 0 0 / 0
Results